OpenStudy (anonymous):

If a circle has an area of 78.5 cm^2, what is its radius? (Show a solution.)

OpenStudy (saifoo.khan):

Use: Area = pi*r^2

OpenStudy (anonymous):

(78.5/(pi))^0.5=r

OpenStudy (anonymous):

sqrt(78.5/pi) = Radius

OpenStudy (anonymous):

\[\large \color{blue}{ r = \sqrt{\frac{Area}{\pi}}}\]